undeadknight26 (undeadknight26):

Solve 3x2 - 12x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work.

undeadknight26 (undeadknight26):

Please help! For part B i have:

undeadknight26 (undeadknight26):

2x^2 - 9x + 7 = 0 (-b +/- √(b^2 - 4ac)) / 2a Where: a = 2 b = 9 c = -7 (-9 +/-sqrt( 9^2 - 4(2)(-7)))/2(2) (-9 +/- sqrt(81 + 56))/4 (-9 +/- sqrt137)/4 x = (-9 + 11.7)/4 x = 2.7/4 x =0.675

OpenStudy (accessdenied):

What other options do you have now, do you have ideas?

undeadknight26 (undeadknight26):

3x^@ - 12x = 0 3x^2 = 12x x^2 = 4x x = 4 This is what i have...is it right?

OpenStudy (accessdenied):

Well it is almost correct. When you divide both sides by x, though, you are making an assumption that x is not equal to 0. However, what happens when you have x=0 in your original equation?

undeadknight26 (undeadknight26):

12x + 0 = 12x so i thought thats how to do it.

OpenStudy (accessdenied):

3x^2 - 12 x = 0 at x=0 3*(0)^2 - 12 * 0 = 0 0 = 0 That is also a solution. :)

undeadknight26 (undeadknight26):

So that means they are both right?

OpenStudy (accessdenied):

Yes. Because you divided both sides by x, which is not allowed when x=0 (division by zero is bad), you removed that information from the equation essentially. So you'd have to test it out afterwards to make sure! A more standard method that takes advantage of the like-factor of x, is by factoring though. 3x^2 - 12x = 0 We could just pull out an x: x (3x - 12) = 0 and use zero-product property. x=0 or 3x - 12 = 0 This will get the same values, without needing to test. :)

undeadknight26 (undeadknight26):

Thank you good sir!

OpenStudy (accessdenied):

You're welcome! :)